a car leaves Rockford, Il at teh rate of 65mph bound for Madison,WI. At the same time, a truck leaves Madison bound for Rockford at the rate of 55 mph. If the cities are 72 miles apart how long will it take for the car and truck to meet?
Answer by lwsshak3(11628)

a car leaves Rockford, Il at teh rate of 65mph bound for Madison,WI. At the same time, a truck leaves Madison bound for Rockford at the rate of 55 mph. If the cities are 72 miles apart how long will it take for the car and truck to meet?
**
let x=distance car traveled
72-x=distance truck traveled
travel time=distance/speed
65/x=55/(72-x)
55x=65*72-65x
120x=65*72
x=65*72/120=39 mi
72-x=33 mi
travel time=39/65=0.6 hr or 36 min
time car and truck met: 36 min